Let f(x) + f(y) = f  (x (1-y2)   +    y (1-x2)  ) . Find the value of        f (4x3 – 3x )  + 3f(x) .

Grade:

1 Answers

Chetan Mandayam Nayakar
11 years ago

f(x)=sin-1(x),f(y)=sin-1(y), and so on...

3f(x)=3sin-1(x)

f(4x3-3x)=f(cos(3cos-1(x)))